03/11/2019 13:29 – sip03.easyipt.co.uk

We are aware of issues with customers routed via sip03.easyipt.co.uk for VoIP calls. We are currently looking at this as a matter of urgency.

UPDATE 01 – 13:39

We have discovered an issue with the Database running on the media gateway and will be performing a reboot. Any active calls will drop.

UPDATE 02 – 13:45

The reboot has completed. However we have lost a number of critical services and are working to restore them.

UPDATE 03 – 13:58

We have been able to recover the services, however concerned about stability and why these services did not automatically start as expected. We are undertaking further reviews and the platform should still be considered as “at risk” until further notice.

UPDATE 04 – 14:47

We have been able to automatically recover a number of services, however we are still seeing some services fail to load on boot. This is something we need to look in too and believe it to be part of a race condition on the server. The media gateway has remained stable and processing calls as expected.

A full review of the media gateway will take place next week to ensure all startup services recover as expected.

03/07/2019 – Web hosting / email retirement and upgrade

At the end of 2019, Windows 2008 server will become EOL (End Of Life) and cease to be supported by Microsoft. In response to this, we will be retiring the platform which we have had in operation since 2011 and migrating accounts to a Debian based operating system well before this.

Unfortunately due to the age of the operating system and the cross vendor platforms there is no direct migration path between systems and will requires changes on both our side and customer applications. (such as emails programs)

The change does bring around a vast number of new features and security options such as SSL (Secure) email which is now becoming industry standard.

All user accounts will need to be migrated by the 30th Aug 2019 when the platform will be turned off. We are currently migrating accounts on a domain by domain basis due to the manual requirement to transfer account data.

Users that have existing accounts setup as IMAP will simply need to delete there accounts when told to and re-create there them with the new server details provided. While we will make every effort to help and backup data, we are limited in what remote assistance we can provide due to unmanaged nature of the product and would strongly suggest users backup there email to local storage prior to making any changes.

Changes our side to DNS will require time to populate around the internet and some disruption is expected on your migration day. Sadly this cant be avoided but we will do all we can to minimize this.

Affected customers will be contacted directly and you are welcome to contact us to expedite your move.

UPDATE 01 – 04/07/2019

In order to ensure data consistency, we have moved to suspend logins to the old platform and would advise anyone who still has access to no make changes until after you have migrated your account as data may be lost.

UPDATE 02 – 10/07/2019

End users will start to see messages about new accounts being created. This is expected as they are being cloned on to the new platform. It also ensures we have up to date contact information for the migration email to follow

UPDATE 03 – 15/07/2019

It has come to our attention that the formatting on the mass email sent out last night has broken. We have uploaded the original email below.

Dear sirs,

Further to our recent advisory on the retirement and renewal of the existing webhosting & email platform. (Details can be found at https://noc.structuredcommunications.co.uk/?p=1387)

Please find below further details and plans on the account migrations that will affect you.


Due to the more modern file structure and use of secure connections on the new platform. Users are advised against simply updating existing email accounts by changing the server details. For ease, users are advised to create a new separate account on there respective email clients and devices using the new server details at the bottom of this email.

Initially, once you create the duplicate account, the contents of the inbox will be out of date. Upon creation of the new account and once you are ready to migrate please visit http://noc.structuredcommunications.co.uk/EasyHTTP/ and complete the form.

Within 24-48 hours the contents of your mailbox will be synchronised to your new account and will begin downloading on your devices. (This may take time if you have a large inbox) Your old account will then stop working and can be deleted from your email client and devices.


Structured Communications will take steps to preserve your email data however is in no way responsible should data loss occur and advise clients to always keep their own data backups.


If you have your email client configured for POP3 based email, then we would advise you to consult your IT personal as the above will result in data loss. You are welcome to contact us if you have questions.


Due to the unmanaged nature of this product and data protection, Structured Communications will be limited in the assistants it can offer end users in setting up and deleting accounts. We can however refer users to one of our IT partners who can offer remote assistance at a charge.


NEW SERVER DETAILS:
USERNAME: Your full email address
PASSWORD: Your existing password

INCOMING SERVER: plesk.easyhttp.co.uk
PORT: 993
ENCRYPTION METHOD: SSL or TLS

OUTGOING SERVER: plesk.easyhttp.co.uk
PORT: 465
ENCRYPTION METHOD: SSL or TLS
OUTBOUND AUTHENTICATION: YES (use same username and password)

Users with a website and custom DNS entries have been mirrored where possible


We understand these changes can be quite confusing and will do all we can do assist and answer any questions however they are necessary.


We will send out reminders for accounts that’s have not migrated before the August 2019 deadline. Should you take no action then your email accounts will stop working and we may lose the ability to sync your inbox data from the date it was initially cloned.


You can contact us on support@structuredcommunications.co.uk or call 0203 301 4000 (even if you want to confirm this email has been sent by ourselves)

UPDATE 04 – 19/07/2019 – User Guides

UPDATE 05 – 01/08/2019 -Helping Users Migrate

We have been working internally to looks at options of migrating current DNS records. This is possible however without the use of SSL due to the IP miss-match between old and new.

In order to assist users in the short term we are moving records requiring little changes if any to email clients. This will allow users to move over before the cut over date automatically, however changes will be required long term as non-SSL connections are turned off.

Domains where no contact has been made will commence from the 05/08/2019. Users should see no interruption.

UPDATE 06 – 30/08/2019 -Extended Deadline

We have, at our own discretion extended the “switch off” date until the 03/09/2019.

09/05/2019 20:50 – Broadband Disruption

We have observed a small number of broadband services suffering from intermittent connection problems between 19:46 and 20:46 this evening.

This issue has been tracked down to one of our wholesale suppliers who suffered a network outage that has since recovered. This has resulted in the affected connections not being able to reach our RADIUS servers for authentication and where simply being terminated on the local BTW RAS with a non route able IP address.

Users who do not have a connection are advised to reboot or power off there router for 20 minutes to recover any stuck sessions.

We apologies for the inconvenience and are awaiting an RFO.

05/04/2019 – SIP VoIP – Outbound Calls

We are aware a small percentage of outbound calls being made on our VoIP network are taking longer than normal to connect. We are aware of this issue and investigating .

UPDATE01 – 09:40

Calls have been re-routed over another carrier while we work to understand what is going on with calls on BTIPX

UPDATE02 – 10:06

Calls are being rejected due to changes made at BT in respect to number formatting. We are in the process of making changes to how we present numbers.

UPDATE03 – 13:00

New number scrips have been designed and put in to operation. Outbound calls are now routing correctly once again and this has been marked as closed.

No changes are required on existing customer systems however we will be changing how we configure new systems

23/02/2019 – 08:00 till 18:00 – Horsham Power – Planned Maintenance

On the 23rd February 2019 starting from 08:00 our UPS vendor Vertiv will be undertaking a planned upgrade to the system adding additional battery capacity to our Horsham facility.

Normally with this type of work the UPS would be placed on “bypass” however the nature of the works require a complete decommission of the system which will require a complete isolation and removal of power to feed “B”

Each cabinet is supplied with a redundant “A” supply. All of our critical services and client facing platforms are dual feed so no outages are expected however should be considered at risk until the completion of the works.

We will be contact clients with co-location who have devices that are not dual redundant.


UPDATE 01 – 08:00

Engineer’s are now on site. Work will start shortly.

UPDATE 02 – 08:30

Work has started and power isolated.

UPDATE 03 – 13.00

Work is now complete and redundancy restored.

01/1/2019 – 22:20 – LNS01 Broadband Gateway

We are aware of a unexpected reload of LNS02 which has occurred again. We are currently dealing with our vendor as to the cause and resolution.

UPDATE 01 – 22:44

Our vendor has responded providing some technical information from the resulting crash and we have some additional logging capability in place. We are unable to advise any further at this stage due to ongoing investigations.

03/01/2019 @ 22:00 – LNS01 Broadband Gateway – Planned Maintenance

We will be taking LNS01 out of service on the 03/01/2019 at 22:00 in order to complete some physical maintenance work for the 04/01/2019. Connections on LNS01 at this time will experience a drop in PPP and reconnect to LNS02.

UPDATE01: 22:48

This work is due to start. Sessions will drop and start to move over to LNS02. If your connection does not come back up within 5 minutes please power cycle your router. If this does not work then please turn off your router for 20 minutes to allow the session to fully close down.

UPDATE02: 23:18

LNS01 has been taking out of service we we have seen 95% of connections move across as expected. Work can now start on this device tomorrow as planned without further impact on services. Updates to follow.

UPDATE03: 12:50

Work has been completed and LNS01 is back in service accepting connections.

12/12/2018 – 21:50 – LNS02 Broadband Gateway

Following on from a unexpected reload of LNS02 and after reviewing the logs we will be upgrading the firmware on this device shortly under an emergency maintiance window.

UPDATE01 – 22:10

The emergency works are now complete on LNS02. This device haas been reloaded and is accepting connections once again. We have provided the logs to the hardware vendor. 

We will need to do the same with LNS01 however, this currently has a large number of active connections and work to move these across the network is needed. This will be done at a later date as we do not believe this device is at risk of the same failure.

We do apologise for the inconvenience caused.

Emergency Work – 23/11/2018 – 21:30 – LNS01 Broadband Gateway

Our network monitoring has flagged high CPU usage on LNS01 that is starting to affect its operation. We are undertaking an emergency reboot to prevent it crashing completely. This will drop active sessions and force them on to other gateways.

UPDATE01 – 21:35
The reboot is complete. Services transferred to redundant gateways as expected. LNS01 is now back in operation. If you do not have service please power down your router for 20 minutes.

We apologise for any inconvenience caused.

22/11/2018 – 09:00 – Outage Vodafone layer2 leased lines

Our network monitoring has alerted us to a small number of metro based layer 2 Vodafone circuits that have no service. This has also had an impact on broadband as well as services in to our Horsham based facility.

This was raised with the relevant teams last night and fibre engineers are already on site working on suspected damaged fibre to our selves and a number of other providers.

We have seen some services restore. Services with backup if affected will be operating on backup and wont be affected.

We will post further updates as they become available.

UPDATE01 – 11:02
Engineers have located the break and this will now progress to splicing.

UPDATE02 – 12:45
Engineers are continuing to work on splicing the damaged cable with spares.

UPDATE03 – 13:00
We are starting to see some services restore.

UPDATE04 – 14:25
As part of ongoing investigation into this issue, Vodafone are sending out additional engineering resources to their data centre.

Vodafone believe the break to be located 43 meters into the fibre handoff, which they are working on to resume full service.

UPDATE05 – 15:30
Engineers at the data centre are continuing to work with Vodafone fibre teams and resolve the second identified fibre break.

UPDATE05 – 16:45
Vodafone fibre engineers continue to examine the length of the fibre for further issues impacting service.

UPDATE06 – 17:55

Vodafone engineers remain on site tracing the underlying infrastructure so that a permanent fix can be put in place for all affected services. We continue to work with the carrier to expedite the resolution with regular communication.

Currently the agreed next stage is for engineers to work to determining the exact location for the fibre fault.

UPDATE07 – 18:45
All services have restored and we are awaiting a RFO. This will be provided on request.